OpenStudy (anonymous):

Simplify the expression -2b-6+3b-2 A b+4 B 5b-8 C b-8 D -b-8 Will medal and fan

OpenStudy (anonymous):

c. b-8

OpenStudy (anonymous):

Look at everything by itself and you have 4 items +3b, -2b, -6, -2 Add similar items together -2b + 3b = b -2- 6 = -8 Put the 2 together and you get b-8 or -8 + b